# Revealing Actual Viscoelastic Relaxation Times in Capillary Breakup

### v1

@@ Abstract
where $G$ is the modulus of the viscoelastic fluid, $D_{1}$ is the initial diameter of the liquid filament, $\gamma$ is the surface tension, and $t_{1}$ is the time of onset of the EC stage [4]. This formula has been adopted in numerous studies to interpret the experimental results of the variation of the filament diameter $D(t)$ during the EC stage. Thus, one obtains an estimate of the apparent relaxation time $\lambda_{e}$ for various kinds of polymer solutions by fitting the exponential decay [15, 12, 30, 31, 25, 26, 27, 29, 32].
@@
However, recent studies have found that pre-stretching of the polymers or filament thinning during the IC/VC stage significantly influences the exponential decay in the subsequent EC stage [14, 17, 18, 23, 32]. Such initial conditions for elasto-capillary thinning can be influenced by multiple parameters, including the initial separation rate of the plates [17] and the initial diameter $D_{0}$ of the liquid bridge [18] in the CaBER method, various nozzle diameters $\ell$ in the dripping method (e.g., experiments indicate $\lambda_{e}\propto\ell^{8/9}$) [23], and different values of the extensibility parameter, $b$, of the polymer in the DoS method [32]. This dependence introduces a strong sensitivity of the extracted relaxation time $\lambda_{e}$ on the initial conditions and inevitable pre-stretching effects that occur in capillary breakup rheometry. To date, there are only a few studies that have indicated the effect of pre-stretching on the apparent relaxation time $\lambda_{e}$ [23, 18, 32], though no studies have identified the actual relaxation time $\lambda$ from capillary breakup measurements, as well as a scaling law of $\lambda_{e}$ versus $\ell$ or $D_{0}$.
@@
The capillary breakup processes considered here include dripping, DoS, and CaBER (Fig. 1a (i, ii, iii)). In all cases, the liquid initially is slowly stretched in a quasi-static manner within devices of characteristic size $\ell$. Under the action of capillary forces driving the Rayleigh-Plateau instability, the liquid rapidly contracts from the initial diameter $D_{0}$ of the VC stage to the initial diameter $D_{1}$ of the EC stage. In our experiments, we used needles with varying outer diameter $\ell\in\{0.3,0.5,0.9,1.8,2.4\}$ mm (corresponding to 30G, 25G, 20G, 15G, and 13G needles) to conduct tests using the dripping and DoS methods. Given that the influence of size of the CaBER method has been tested in Ref. [18], we also used their data of the PEOvis solution (see [33], §I and Fig. S1) for our analysis.
Three kinds of polymer solutions were employed in our experiments: PIB-PB (polyisobutylene dissolved in a polybutene-mineral oil mixed solvent), PS-DOP (polystyrene dissolved in the organic solvent dioctyl phthalate), and PEO-PEG (polyethylene oxide dissolved in a polyethylene glycol-water solution), whereas PIB-PB-0.3 and PIB-PB-0.02 denote different PIB concentrations; PEO-PEG-8M and PEO-PEG-1M identify different molecular weights of PEO. These latter fluids can be regarded as Boger fluids with constant viscosity (Fig. S2 [33]), ensuring that additional effects from viscosity variations with extension rate are eliminated. Detailed preparation procedures and physical property measurements, such as viscosity $\eta_{0}$, surface tension $\gamma$, storage modulus $G^{\prime}$, and loss modulus $G^{\prime\prime}$, are provided in [33] §II and summarized in Table S1.
@@
Figure 2: Dependence of apparent relaxation time $\lambda_{e}$ on apparatus size $\ell$ for different Boger fluids using specific methods: PEO-PEG, PS-DOP, PIB-PB in this work by dripping and/or DoS, and PEOvis (Ref. [18]) by CaBER.
@@
$\frac{1}{Ec}\frac{1}{a}=\frac{3\beta}{1-\beta}E+f(A_{zz}-A_{rr})$ (2a)
@@
where we have introduced dimensionless time $\tau=t/\lambda$, elasto-capillary number $Ec=\eta_{p}D_{0}/2\lambda\gamma$, solvent viscosity ratio $\beta=\eta_{s}/\eta_{0}$, dimensionless (time-varying) extensional strain rate $E=-2\lambda D^{-1}\frac{dD}{dt}$, the conformation tensor $\bm{A}$ with $rr$ and $zz$ components, and the finite extensibility function $f=(b-3)/(b-A_{zz}-2A_{rr})$. By comparing simplified analytical solutions (see [33] §IV-B1, B2), approximate solutions ([33] §IV-B3), and an analysis of the full model ([33] §IV-C1) under different assumptions, we demonstrated that the evolution is sensitive to the initial conformation $\bm{A}(t_{1})$ for the EC stage. Only by solving the full model numerically can the variations in the VC, EC, and VC-EC transitions be captured accurately (see [33] §IV).
@@
For a specific viscoelastic liquid, as properties $\eta_{p}$, $\beta$ and $\gamma$ are measurable and size $D_{0}$ is controlled, we can assume a true relaxation time $\lambda$ for determining $Ec$ in Eq. (2a), and explore a wide range of $b$ values in the numerical calculations. Taking the PEO-PEG solution as an example, numerical predictions of the relative ratio $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ (Fig. 3a, c, e) and $\lambda_{e}/\lambda$ (Fig. 3b, d, f) for assumed $\lambda\in\{5,6,7\}\,\mathrm{ms}$ and $b\in\{10^{4},2\times 10^{4},4\times 10^{4},8\times 10^{4}\}$ are compared with experimental results (symbols in Fig. 3). The trend of $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ for different $D_{0}$ values provides a significant indicator for comparison with model predictions because both $\lambda_{e}/\lambda$ and $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ are uniquely determined for fixed $E_{c}$ and $\beta$.
@@
Figure 4: Dimensionless map of the relaxation time ratio $\lambda_{e}/\lambda$ versus the measurable elasto-capillary number $Ec_{e}=D_{0}\eta_{p}/(2\gamma\lambda_{e})$ for various unknown extensibility values $b\in[10^{2},10^{6}]$. (a) 3D maps for different $\beta$ of a specific solution. (b) Top view of 3D maps. (c) Collapsed 2D map for easy comparison. The symbols represent experimental results (red: PEO-PEG-8M, purple: PEO-PEG-1M, green: PIB-PB-0.3, black: PIB-PB-0.02, blue: PS-DOP) and Ref. [18] (gray: PEOvis). Solid lines are obtained from fitting, with deviation ranges enclosed by dash-dot lines, suggesting the estimations.
Next, we analyzed the experimental data presented in Fig. 2 and constructed a dimensionless phase diagram (note the logarithmic axes) of $\lambda_{e}/\lambda$ as a function of $Ec_{e}$, $b$ and $\beta$ (Fig. 4a), where $Ec_{e}=D_{0}\eta_{p}/(2\gamma\lambda_{e})$, with each variable being directly measurable from experiments. Although different $\beta$ values affect the map, particularly for small $b$, indicated in Fig. 4b, we can construct the phase diagram with all experimental data as shown in Fig. 4c; uncertainties in all measurable parameters were considered to determine the true relaxation time $\lambda$ and the range of extensibility $b$ for each viscoelastic fluid tested.
@@
From the phase diagram (Fig. 4), two distinct regions can be identified. When $\log_{10}(Ec_{e}b)\gtrsim 1.5$, the experimentally measured $\lambda_{e}$ becomes independent of the experimental size, i.e. $\lambda_{e}\sim\ell^{0}$. However, this does not imply that the true relaxation time $\lambda$ is obtained. In this region, since $b$ cannot be experimentally fitted, an additional evaluation of $b$ is necessary to determine the specific value of $\lambda_{e}/\lambda$. In the region where $\log_{10}(\lambda_{e}/\lambda)\lesssim-0.6$, the slope of $\log_{10}(\lambda_{e}/\lambda)$ approaches infinity, indicating that $\lambda_{e}\sim D_{0}$. This is consistent with the scaling law $\lambda_{e}\sim\ell$ observed in Fig. 2 when $\ell(D_{0})$ becomes sufficiently small. The $\lambda_{e}\sim\ell^{0.5}$ scaling law observed in Fig. 2 represents an empirical scaling behavior characterizing the transition between the two regions.
@@
Thus, some predictions can be made. For fluids with relatively low surface tension and strong viscoelasticity, such as PIB-PB-0.3 and PS-DOP, which exhibit $\log_{10}(Ec_{e})\gtrsim-1.5$ during capillary breakup, a size-independent exponential decay in the EC stage may occur. This implies that a nearly constant apparent relaxation time can be observed. In contrast, fluids with relatively high surface tension and low viscoelasticity, such as water-based dilute solutions like PEO-PEG and PEOvis, exhibit significant size dependence of the apparent relaxation time. The extensibility $b$ determines the value of the critical $Ec_{e}$ at which this effect occurs, with smaller $b$ making this influence more readily triggered. Additionally, we speculate that in micro-scale capillary breakup processes (e.g., within biological tissues), where $Ec_{e}$ is extremely small, the linear dependence of $\lambda_{e}\sim D_{0}$ is likely to occur.
@@
In summary, our study reveals, through experiments and numerical solutions, how inevitable pre-stretching during the visco-capillary stage of capillary breakup influences the exponential decay in the elasto-capillary stage and the apparent relaxation time $\lambda_{e}$ obtained using classical formulations for viscoelastic fluids. Through a theoretical model, we demonstrate that the apparent relaxation time is always smaller than the actual relaxation time, i.e., $\lambda_{e}<\lambda$. We further identify that the size dependence of $\lambda_{e}$ is governed by the extensibility $b$, solvent viscosity ratio $\beta$, and the dimensionless number $Ec_{e}=D_{0}\eta_{p}/(2\gamma\lambda_{e})$. Large $b$ and $Ec_{e}$ result in size-independent behavior, while under small $Ec_{e}$ conditions, the scaling $\lambda_{e}\sim D_{0}\sim\ell$ emerges. We present a methodology for determining the actual relaxation time $\lambda$ by synthesizing a series of apparent relaxation times $\lambda_{e}$ measured under various size conditions. These findings provide a new perspective on accurately measuring viscoelastic properties using capillary breakup rheometry and offer insights into the size effects in the capillary breakup dynamics of viscoelastic fluids. Looking forward, several factors need to be considered to address the measurement of more complex fluids, such as those with shear-thinning viscosity, fluids where the IC stage dominates instead of the VC stage.

### v2

@@ Abstract
where $G$ is the modulus of the viscoelastic fluid, $D_{1}$ is the initial diameter of the liquid filament, $\gamma$ is the surface tension, and $t_{1}$ is the time of onset of the EC stage [4]. This formula has been adopted in numerous studies to interpret the experimental results of the variation of the filament diameter $D(t)$ during the EC stage. Thus, one obtains an estimate of the apparent relaxation time $\lambda_{e}$ for various kinds of polymer solutions by fitting the exponential decay [15, 12, 30, 31, 25, 26, 27, 29, 32, 33].
@@
However, existing studies have found that pre-stretching of the polymers or filament thinning during the IC/VC stage significantly influences the exponential decay in the subsequent EC stage [14, 17, 18, 23, 32]. The sensitivity of $\lambda_{e}$ to these effects can be attributed to several experimental parameters, including the initial separation rate of the plates [17], the initial diameter $D_{0}$ of the liquid bridge [18] in the CaBER method, nozzle diameter $\ell$ in the dripping method (where $\lambda_{e}\propto\ell^{8/9}$ has been observed) [23], and the polymer extensibility parameter $b$ in the DoS method [32]. This parameter dependence introduces strong sensitivity of the extracted relaxation time $\lambda_{e}$ to initial conditions and inevitable pre-stretching effects in capillary breakup rheometry. While several studies have reported such dependence [23, 18, 32], no work has yet directly identified the actual relaxation time $\lambda$ from breakup measurements, nor established a general scaling law relating $\lambda_{e}$ to $\ell$ or $D_{0}$. In contrast, a recent study [33] reported no systematic dependence of $\lambda_{e}$ on CaBER plate size across four polymer solutions with varying concentration and molecular weight. This highlights the need for a unified framework to account for the size-dependent variability of $\lambda_{e}$ observed in different experimental systems.
@@
The capillary breakup processes considered here include dripping, DoS, and CaBER (Fig. 1a (i, ii, iii)). In all cases, the liquid initially is slowly stretched in a quasi-static manner within devices of characteristic size $\ell$. Under the action of capillary forces driving the Rayleigh-Plateau instability, the liquid rapidly contracts from the initial diameter $D_{0}$ of the VC stage to the initial diameter $D_{1}$ of the EC stage. In our experiments, we used needles with varying outer diameter $\ell\in\{0.3,0.5,0.9,1.8,2.4\}$ mm (corresponding to 30G, 25G, 20G, 15G, and 13G needles) to conduct tests using the dripping and DoS methods. Given that the influence of the size of the CaBER method has been tested in Ref. [18, 33], we also used their data of the PEOvis and PS solution (see [34], §I and Fig. S1) for our analysis.
Three kinds of polymer solutions were employed in our experiments: PIB-PB (polyisobutylene dissolved in a polybutene-mineral oil mixed solvent), PS-DOP (polystyrene dissolved in the organic solvent dioctyl phthalate), and PEO-PEG (polyethylene oxide dissolved in a polyethylene glycol-water solution), whereas PIB-PB-0.3 and PIB-PB-0.02 denote different PIB concentrations; PEO-PEG-8M and PEO-PEG-1M identify different molecular weights of PEO. These latter fluids can be regarded as Boger fluids with constant viscosity (Fig. S2 [34]), ensuring that additional effects from viscosity variations with extension rate are eliminated. Detailed preparation procedures and physical property measurements, such as viscosity $\eta_{0}$, surface tension $\gamma$, storage modulus $G^{\prime}$, and loss modulus $G^{\prime\prime}$, are provided in [34] §II and summarized in Table S1. The polymer concentrations $c$ used in this study fall within the dilute regime, with $c^{*}>c\gtrsim 0.1c^{*}$, and are all significantly higher than the minimal concentrations required for EC stage [35, 36]. These include $c_{\mathrm{min}}\sim c^{*}/b$ and $c_{\mathrm{low}}\sim c^{*}/b^{0.75}$, which represent the thresholds for the emergence of the EC regime [35] and for visibly observable exponential thinning [36], respectively.
@@
Figure 2: Dependence of apparent relaxation time $\lambda_{e}$ on apparatus size $\ell$ for different Boger fluids using specific methods: PEO-PEG, PS-DOP, PIB-PB in this work by dripping and/or DoS, PEOvis (Ref. [18]) and PS16/PS7 (Ref. [33]) by CaBER.
@@
$\frac{Ec}{a}=\frac{3\beta}{1-\beta}E+f(A_{zz}-A_{rr})$ (2a)
@@
where we have introduced dimensionless time $\tau=t/\lambda$, elasto-capillary number $Ec=2\lambda\gamma/\eta_{p}D_{0}$, solvent viscosity ratio $\beta=\eta_{s}/\eta_{0}$, dimensionless (time-varying) extensional strain rate $E=-2\lambda D^{-1}\frac{dD}{dt}$, the conformation tensor $\bm{A}$ with $rr$ and $zz$ components, and the finite extensibility function $f=(b-3)/(b-A_{zz}-2A_{rr})$. By comparing simplified analytical solutions (see [34] §IV-B1, B2), approximate solutions ([34] §IV-B3), and an analysis of the full model ([34] §IV-C1) under different assumptions, we demonstrated that the evolution is sensitive to the initial conformation $\bm{A}(t_{1})$ for the EC stage. Only by solving the full model numerically can the variations in the VC, EC, and VC-EC transitions be captured accurately (see [34] §IV).
@@
For a specific viscoelastic liquid, as properties $\eta_{p}$, $\beta$ and $\gamma$ are measurable and size $D_{0}$ is controlled, we can assume a true relaxation time $\lambda$ for determining $Ec$ in Eq. (2a), and explore a wide range of $b$ values in the numerical calculations. Taking the PEO-PEG solution as an example, numerical predictions of the relative ratio $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ (Fig. 3a, c, e) and $\lambda_{e}/\lambda$ (Fig. 3b, d, f) for assumed $\lambda\in\{5,6,7\}\,\mathrm{ms}$ and $b\in\{10^{4},2\times 10^{4},4\times 10^{4},8\times 10^{4}\}$ are compared with experimental results (symbols in Fig. 3). The trend of $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ for different $D_{0}$ values provides a significant indicator for comparison with model predictions because both $\lambda_{e}/\lambda$ and $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ are uniquely determined for fixed $Ec$ and $\beta$.
@@
Figure 4: Dimensionless map of the relaxation time ratio $\lambda_{e}/\lambda$ versus the measurable elasto-capillary number $Ec_{e}=2\gamma\lambda_{e}/(D_{0}\eta_{p})$ for various unknown extensibility values $b\in[10^{2},10^{6}]$. (a) 3D maps for different $\beta$ of a specific solution. (b) Top view of 3D maps. (c) Collapsed 2D map for easy comparison. The symbols represent experimental results (red: PEO-PEG-8M, purple: PEO-PEG-1M, green: PIB-PB-0.3, black: PIB-PB-0.02, blue: PS-DOP) from current work, Ref. [18] (gray: PEOvis), and Ref. [33] (light/dark orange: PS16-50/100, light/dark brown: PS7-200/500). Solid lines are obtained from fitting, with deviation ranges enclosed by dash-dot lines, suggesting the estimations.
Next, we analyzed the experimental data presented in Fig. 2 and constructed a dimensionless phase diagram (note the logarithmic axes) of $\lambda_{e}/\lambda$ as a function of $Ec_{e}$, $b$ and $\beta$ (Fig. 4a), where $Ec_{e}=2\gamma\lambda_{e}/(D_{0}\eta_{p})$, with each variable being directly measurable from experiments. Although different $\beta$ values affect the map, particularly for small $b$, indicated in Fig. 4b, we can construct the phase diagram with all experimental data as shown in Fig. 4c; uncertainties in all measurable parameters were considered to determine the true relaxation time $\lambda$ and the range of extensibility $b$ for each viscoelastic fluid tested.
@@
From the phase diagram (Fig. 4), two distinct regions can be identified. When $\log_{10}(Ec_{e}^{-1}b)\gtrsim 1.5$, the experimentally measured $\lambda_{e}$ becomes independent of the experimental size, i.e. $\lambda_{e}\sim\ell^{0}$. However, this does not imply that the true relaxation time $\lambda$ is obtained. In this region, since $b$ cannot be experimentally fitted, an additional evaluation of $b$ is necessary to determine the specific value of $\lambda_{e}/\lambda$. In the region where $\log_{10}(\lambda_{e}/\lambda)\lesssim-0.6$, the slope of $\log_{10}(\lambda_{e}/\lambda)$ approaches infinity, indicating that $\lambda_{e}\sim D_{0}$. This is consistent with the scaling law $\lambda_{e}\sim\ell$ observed in Fig. 2 when $\ell(D_{0})$ becomes sufficiently small. The $\lambda_{e}\sim\ell^{0.5}$ scaling law observed in Fig. 2 represents an empirical scaling behavior characterizing the transition between the two regions.
@@
Thus, some predictions can be made. For fluids with relatively low surface tension and strong viscoelasticity, such as PIB-PB-0.3 and PS-DOP, which exhibit $\log_{10}(Ec_{e})\lesssim 1.5$ during capillary breakup, a size-independent exponential decay in the EC stage may occur. This implies that a nearly constant apparent relaxation time can be observed. In contrast, fluids with relatively high surface tension and low viscoelasticity, such as water-based dilute solutions like PEO-PEG and PEOvis, exhibit significant size dependence of the apparent relaxation time. The extensibility $b$ determines the value of the critical $Ec_{e}$ at which this effect occurs, with smaller $b$ making this influence more readily triggered. Additionally, we speculate that in micro-scale capillary breakup processes (e.g., within biological tissues), where $Ec_{e}$ is extremely big, the linear dependence of $\lambda_{e}\sim D_{0}\sim\ell$ is likely to occur.
@@
In summary, our study reveals, through experiments and numerical solutions, how inevitable pre-stretching during the visco-capillary stage of capillary breakup influences the exponential decay in the elasto-capillary stage and the apparent relaxation time $\lambda_{e}$ obtained using classical formulations for viscoelastic fluids. Through a theoretical model, we demonstrate that the apparent relaxation time is always smaller than the actual relaxation time, i.e., $\lambda_{e}<\lambda$. We further identify that the size dependence of $\lambda_{e}$ is governed by the extensibility $b$, solvent viscosity ratio $\beta$, and the dimensionless number $Ec_{e}=2\gamma\lambda_{e}/(D_{0}\eta_{p})$. Large $b$ and small $Ec_{e}$ result in size-independent behavior, while under big $Ec_{e}$ conditions, the scaling $\lambda_{e}\sim D_{0}\sim\ell$ emerges. We present a methodology for determining the actual relaxation time $\lambda$ by synthesizing a series of apparent relaxation times $\lambda_{e}$ measured under various size conditions. These findings provide a new perspective on accurately measuring viscoelastic properties using capillary breakup rheometry and offer insights into the size effects in the capillary breakup dynamics of viscoelastic fluids. Looking forward, several factors need to be considered to address the measurement of more complex fluids, such as those with shear-thinning viscosity, fluids where the IC stage dominates instead of the VC stage.